OpenStudy (anonymous):

a company uses 2 vans to transport workers from a free parking lot to the workplace between 7 am to 9am. one vans has 9 more seats the other. the smaller van takes two trips every morning while the larger van takes only one trip. the 2 vans can transaport 69 people, maximum. let x be the seats in the small van and y the seats in the large van. how many seats does the larger van have?

OpenStudy (anonymous):

y=x+9 y+2x=69 x+9+2x=69 x=20 y=29

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome